# Robust Matrix Completion for Discrete Rating-Scale Data

### v1

@@ Abstract
Matrix completion has gained considerable interest in recent years. The goal of matrix completion is to predict the unknown entries of a partially observed matrix using its known entries. Although common applications feature discrete rating-scale data, such as user-product rating matrices in recommender systems or surveys in the social and behavioral sciences, methods for matrix completion are almost always designed for and studied in the context of continuous data. Furthermore, only a small subset of the literature considers matrix completion in the presence of corrupted observations despite their common occurrence in practice. Examples include attacks on recommender systems (i.e., malicious users deliberately manipulating ratings to influence the recommender system to their advantage), or careless respondents in surveys (i.e., respondents providing answers irrespective of what the survey asks of them due to a lack of attention). We introduce a matrix completion algorithm that is tailored towards the discrete nature of rating-scale data and robust to the presence of corrupted observations. In addition, we investigate the performance of the proposed method and its competitors with discrete rating-scale (rather than continuous) data as well as under various missing data mechanisms and types of corrupted observations.
Keywords ADMM algorithm, outliers, latent low-rank regularization, recommender systems, surveys
@@
This work is supported by a grants of the Dutch Research Council (NWO), research program Vidi, grant numbers VI.Vidi.195.141 (Andreas Alfons) and VI.Vidi.211.032 (Ines Wilms).
@@ 1 Introduction
Matrix completion aims to predict the unknown entries of a partially observed matrix using its known elements by minimizing a criterion such as the mean squared error, typically subject to a low-rank constraint. While early work on matrix completion dates back to, among others, Achlioptas & McSherry, (2001), Srebro et al., (2004), and Rennie & Srebro, (2005), it has been actively studied since Candès & Recht, (2009) and Candès & Tao, (2010). For overviews, we refer to, e.g., Mazumder et al., (2010), Hastie et al., (2015), Davenport & Romberg, (2016), Fithian & Mazumder, (2018), and Nguyen et al., (2019). Matrix completion has been successfully applied in a variety of fields, ranging from the famous Netflix Prize competition (e.g., Rennie & Srebro,, 2005; Bennett & Lanning,, 2007; see LeBlanc et al.,, 2024, for a recent review), over image manifolds in computer vision and pattern recognition (e.g., Weinberger & Saul,, 2006) or sensor positioning in sensor network localization (e.g., So & Ye,, 2007), to modeling disease progressions in medical practice and clinical research (e.g., Kidziński & Hastie,, 2023).
@@
A multitude of methods for matrix completion exist, including weighted low-rank matrix approximation through singular value decomposition (e.g., Eckart & Young,, 1936; Srebro & Jaakkola,, 2003); matrix factorization (e.g., Srebro et al.,, 2004; Rennie & Srebro,, 2005; or Schiavon et al.,, 2024, for recent computational advances); rank minimization, among which nuclear norm minimization (e.g., Candès & Recht,, 2009; Candès & Tao,, 2010; Mazumder et al.,, 2010), singular value thresholding (e.g., Cai et al.,, 2010; Chatterjee,, 2015; Lei & Zhou,, 2019), or iteratively reweighted least squares minimization (e.g., Mohan & Fazel,, 2012); or combinations of the former (e.g., Hastie et al.,, 2015; Cho et al.,, 2019). In this paper, we contribute to the literature by presenting a matrix completion procedure called robust discrete matrix completion (RDMC) that is tailored towards the discrete nature of incomplete—oftentimes missing not at random—rating-scale data, while being robust to the presence of corrupted observations.
@@
First, the vast majority of existing methods for matrix completion are formulated over the real number domain and produce continuous predictions, despite important applications featuring (discrete and bounded) rating-scale data. Exceptions that include discreteness or box constraints can be found in Huang et al., (2013), Huo et al., (2016), Nguyen et al., (2018), Tatsukawa & Tanaka, (2018), Iimori et al., (2020), or Bertsimas & Li, (2023). One major application, and arguably one of the key drivers for the interest in matrix completion, are recommender systems: users rate items (such as movies, songs, or any other products) according to their liking, for instance from one star to five stars, with predictions being used to recommend new items to users. Another relevant application, although rarely considered in the literature on (low-rank) matrix completion, are surveys in the social and behavioral sciences: researchers measure each latent construct of interest (such as personality traits) by multiple rating-scale items, with missing values commonly occurring due to nonresponse in parts of the survey. Although the observed ratings can be interpreted as discrete measurements of a latent continuous sentiment, the underlying continuous distributions are not identified without additional assumptions (e.g., Bond & Lang,, 2019), making it impossible to compare predictions across different items. The proposed procedure therefore restricts the predictions to the given rating scale via a discreteness constraint.
Second, only a relatively small subset of the literature considers matrix completion in the presence of outliers or corrupted/manipulated observations. Existing approaches are commonly based on robust nuclear norm minimization (e.g., Chen et al.,, 2013; Huang et al.,, 2013; Nie et al.,, 2015; Cambier & Absil,, 2016; Elsener & van de Geer,, 2018; Shang & Kong,, 2021) or robust matrix factorization (e.g., Zhao et al.,, 2016; Zhang et al.,, 2017; Ruppel et al.,, 2020; Tang & Guan,, 2020; Nie et al.,, 2021; Wang & Fan,, 2024). Corrupted observations frequently occur in practice, and despite the bounded nature of rating-scale data, they can have a substantial impact on predictions. In recommender systems applications, for instance, malicious agents may launch so-called attacks with fake user profiles to deliberately manipulate recommendations to their advantage (e.g., Van Roy & Yan,, 2010; see also Gunes et al.,, 2014, and Si & Li,, 2020, for overviews). In survey applications, so-called careless responding (i.e., responses not based on item content due to inattention or misunderstanding; e.g., Ward & Meade,, 2023) or bot responding (e.g., Storozuk et al.,, 2020) are common occurrences in online surveys, and a prevalence as low as 5% can invalidate research findings (Credé,, 2010; Arias et al.,, 2020; Welz et al.,, 2024). Both of these examples highlight the need to protect against corrupted observations in matrix completion for discrete rating-scale data, which the proposed procedure incorporates through a robust loss function.
@@
Third, much of the literature does not investigate the effect of a missing not at random (MNAR) mechanism on matrix completion procedures, with a notable exception being Choi & Yuan, (2024). Yet missing values are frequently MNAR in practice. In recommender systems applications, users may predominantly rate items that they have consumed or purchased and thus expected to like, intrinsically linking the probability of missingess to their preferences. In survey applications, participants may abandon the survey yielding higher probability of missingness for later survey items, or they may not respond to items they find, e.g., confusing or inappropriate. Hence, we study the performance of the proposed method under realistic MNAR settings via extensive simulation studies in addition to empirical applications.
@@
The remainder of the paper is structured as follows. Section 2 presents the proposed procedure for robust matrix completion with discrete rating-scale data. In Section 3, the performance of our procedure is evaluated via simulation studies for recommender systems and rating-scale surveys, while two empirical applications are discussed in Section 4. The final Section 5 concludes the paper.
@@ 2 A robust procedure for discrete matrix completion
We start by formulating the regularized optimization problem in Section 2.1 and presenting the corresponding algorithm in Section 2.2. Section 2.3 provides further details on centering the columns of the observed data matrix and initializing the algorithm, while Section 2.4 introduces different choices of robust loss functions. Finally, the selection of the regularization parameter is discussed in Section 2.5. Our implementation in the add-on package rdmc (Alfons & Archimbaud,, 2024) for the statistical computing environment R (R Core Team,, 2024) is freely available from https://github.com/aalfons/rdmc. The main part of the code is thereby written in C++ to improve computational efficiency.
@@ 2.1 Problem formulation
Suppose that we observe an incomplete rating matrix $\bm{X}$ of $n$ rows (representing individuals providing the ratings) and $p$ columns (representing the rated items) with elements $X_{ij}\in\mathcal{C}_{j}=\{c_{1}^{j},\dots,c_{K}^{j}\}\subset\mathbb{R}$, where $K$ denotes the number of discrete rating categories and $c_{1}^{j}<\dots<c_{K}^{j}$ for $j=1,\dots,p$. Note that the values of the rating categories may vary between columns, which is related to column centering as further discussed in Section 2.3. Let $\Omega$ denote the index set of observed entries in ${\bf X}$ and define the projection $P_{\Omega}$ to be the $(n\times p)$-dimensional matrix whose elements are given by
$\left(P_{\Omega}({\bf X})\right)_{ij}=\begin{cases}X_{ij}&\text{ if }(i,j)\in\Omega,\\ 0&\text{ otherwise.}\end{cases}$
@@
In order to preserve the discrete nature of the ratings, one could add a discreteness constraint to (1) (cf. Huang et al.,, 2013). However, requiring the elements of $\bm{L}$ to be discrete may not yield a solution, since such a discreteness constraint and a rank constraint are unlikely to be fulfilled simultaneously. To resolve this issue, an ancillary continuous matrix $\bm{Z}$ can be introduced such that the discreteness constraint operates on $\bm{L}$ and the nuclear norm constraint on $\bm{Z}$, while ensuring that $\bm{L}$ and $\bm{Z}$ remain as close as possible. In addition, the squared Frobenius norm in (1) does not protect against corrupted observations (such as fake profiles in recommender systems or careless respondents in surveys). For increased protection, we replace it with a (pseudo-)norm $\|\bm{Y}\|_{\rho}=\sum_{i,j}\rho(Y_{ij})$ based on a robust loss function $\rho$ (cf. Tang & Guan,, 2020). Putting all of this together, we obtain the following optimization problem in augmented Lagrangian form:
@@
where $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product, $\bm{\Theta}$ is a multiplier adjusting for the discrepancy between $\bm{L}$ and $\bm{Z}$, and $\mu$ is an additional regularization parameter. It should be noted that (2) generalizes the optimization problem formulated in Huang et al., (2013).
@@ 2.2 Algorithm
$L_{ij}=\begin{cases}{\displaystyle\operatorname*{arg\,min}_{c_{k}\in\mathcal{C}_{j}}}\ \rho(c_{k}-X_{ij})+\frac{\mu}{2}(c_{k}-Z_{ij}+\frac{1}{\mu}\Theta_{ij})^{2}&\text{ for }(i,j)\in\Omega,\\ {\displaystyle\operatorname*{arg\,min}_{c_{k}\in\mathcal{C}_{j}}}\ (c_{k}-Z_{ij}+\frac{1}{\mu}\Theta_{ij})^{2}&\text{ for }(i,j)\notin\Omega.\end{cases}$
Third, update the discrepancy parameter $\bm{\Theta}\leftarrow\bm{\Theta}+\mu(\bm{L}-\bm{Z})$ and the regularization parameter $\mu\leftarrow\delta\mu$ with $\delta>1$. Hence $\mu$ grows exponentially in the number of iterations to speed up convergence.
Following Huang et al., (2013) and Tang & Guan, (2020), we initialize $\mu=0.1$ and set $\delta=1.05$. As convergence criterion, we take the relative change in the objective function from (2) falling below a given threshold, which we set to $10^{-4}$.

### v2

@@ Abstract
Recommender systems are essential tools in the digital landscape for connecting users with content that more closely aligns with their preferences. Matrix completion is a widely used statistical framework for such systems, aiming to predict a user’s preferences for items they have not yet rated by leveraging the observed ratings in a partially filled user–item rating matrix. Realistic applications of matrix completion in recommender systems must address several challenges that are too often neglected: (i) the discrete nature of rating-scale data, (ii) the presence of malicious users who manipulate the system to their advantage through the creation of fake profiles, and (iii) missing-not-at-random patterns, where users are more likely to rate items they expect to enjoy. Our goal in this paper is twofold. First, we propose a novel matrix completion method, robust discrete matrix completion (RDMC), designed specifically to handle the discrete nature of sparse rating-scale data and to remain reliable in the presence of adversarial manipulation. We evaluate RDMC through carefully designed experiments and realistic case studies. Our work therefore, secondly, offers a statistically-sound blueprint for future studies on how to evaluate matrix completion methods for recommender systems under realistic scenarios.
Keywords adversarial attacks, imputation, latent low-rank regularization, missing data, outliers, recommendation engines
@@
This work is supported by grants of the Dutch Research Council (NWO), research program Vidi, grant numbers VI.Vidi.195.141 (Andreas Alfons) and VI.Vidi.211.032 (Ines Wilms).
@@ 1 Introduction
Recommender systems play a critical role in modern digital platforms, particularly in online retail and advertising, where they help users discover relevant products efficiently and drive substantial commercial value. Improvements in recommendation accuracy, even marginally, can yield significant financial gains in industries such as e-commerce and computational advertising (LeBlanc et al.,, 2024). However, malicious agents may launch so-called attacks with fake user profiles to deliberately manipulate recommendations to their advantage (e.g., Van Roy & Yan,, 2010; see also Gunes et al.,, 2014, and Si & Li,, 2020, for overviews). For instance, already in the early 2000s, platforms like Amazon and eBay have reported repeated attempts to subvert their recommender systems through deceptive practices such as fake reviews and purchased ratings (Lam & Riedl,, 2004). Today, these concerns remain more relevant than ever, as highlighted by Adamopoulos, (2024), and they are drawing increasing attention from regulatory authorities such as the Federal Trade Commission, (2019), the Competition and Markets Authority, (2021), or the European Commission, (2022). For instance, the Federal Trade Commission, (2024) recently introduced new measures and regulations aimed at protecting consumers, which include urging online platforms to proactively detect and address fake profiles and suspicious behavioral patterns—such as those generated by artificial intelligence. This paper introduces a novel matrix completion method that can withstand adversarial manipulation in recommender systems; our experiments offer a statistically-sound yet realistic blueprint for future studies to evaluate performance.
From a statistical perspective, recommender systems are often framed as matrix completion problems for a sparse user-item rating matrix (e.g., ranging from one star to five stars), with predictions being used to recommend new items to users. A prominent example is the famous Netflix Prize competition (Bennett & Lanning,, 2007). While early work on matrix completion dates back to, among others, Achlioptas & McSherry, (2001), Srebro et al., (2004), and Rennie & Srebro, (2005), it has been actively studied since Candès & Recht, (2009) and Candès & Tao, (2010). For overviews, we refer to, e.g., Davenport & Romberg, (2016), Nguyen et al., (2019), and LeBlanc et al., (2024).
@@
A multitude of methods for matrix completion exist, including weighted low-rank matrix approximation through singular value decomposition (e.g., Srebro & Jaakkola,, 2003); matrix factorization (e.g., Srebro et al.,, 2004; Rennie & Srebro,, 2005; or Schiavon et al.,, 2024, for recent computational advances); rank minimization, among which nuclear norm minimization (e.g., Candès & Recht,, 2009; Candès & Tao,, 2010; Mazumder et al.,, 2010), singular value thresholding (e.g., Cai et al.,, 2010; Chatterjee,, 2015; Lei & Zhou,, 2019), iteratively reweighted least squares minimization (e.g., Mohan & Fazel,, 2012); combinations of the former (e.g., Hastie et al.,, 2015; Cho et al.,, 2019); or dynamic approaches over time (e.g., Chen et al.,, 2024).
We contribute to the literature, first methodologically, by presenting a novel matrix completion procedure called robust discrete matrix completion (RDMC) that is tailored towards the discrete nature of incomplete rating-scale data while being robust to adversarial manipulation. We assess RDMC against the widely-used procedure Soft-Impute (Mazumder et al.,, 2010; Hastie et al.,, 2015), as well as a variant that we adjusted with a post-hoc discretization step to fit the discrete nature of rating-scale data. We study these methods in empirical case studies and extensive simulations. These experiments are carefully designed to be more realistic than previous studies, thereby providing an important second conceptual contribution to the literature, as our experiments may serve as realistic yet statistically-sound blueprints for future studies in the context of matrix completion and recommender systems to further boost transparency, reproducibility and generalizability. With our contributions, we in particular address the following issues.
@@
First, the vast majority of existing methods for matrix completion are formulated over the real number domain and produce continuous predictions, despite recommender systems applications featuring (discrete and bounded) rating-scale data. Exceptions that include discreteness or box constraints can be found in Huang et al., (2013), Huo et al., (2016), Nguyen et al., (2018), Tatsukawa & Tanaka, (2018), Iimori et al., (2020), or Bertsimas & Li, (2023). Although the observed ratings may be interpreted as discrete measurements of a latent continuous sentiment, the underlying continuous distributions are not identified without additional assumptions (cf. Bond & Lang,, 2019), invalidating comparisons of predictions across different items. The proposed procedure RDMC therefore restricts the predictions to the given rating scale via a discreteness constraint.
Second, only a relatively small subset of the literature on matrix completion considers the presence of outliers or corrupted/manipulated observations, such as fake profiles in recommender systems. From a perspective of robust statistics, there are primarily two approaches for handling outliers to avoid biased analyses. One approach involves explicitly identifying and removing outliers, while the generally favored approach focuses on robust methods that can accommodate the presence of outliers without compromising performance (see, e.g., Avella Medina & Ronchetti,, 2015, for an overview). In the context of matrix completion and recommender systems, an example of the first strategy is given in Lee & Zhu, (2012), while existing procedures for the latter are commonly based on robust nuclear norm minimization (e.g., Chen et al.,, 2013; Huang et al.,, 2013; Nie et al.,, 2015; Cambier & Absil,, 2016; Elsener & van de Geer,, 2018; Shang & Kong,, 2021) or robust matrix factorization (e.g., Zhao et al.,, 2016; Zhang et al.,, 2017; Ruppel et al.,, 2020; Tang & Guan,, 2020; Wang & Fan,, 2025). Among these, only Huang et al., (2013) address discrete data. Moreover, Adamopoulos, (2024) cautions against the prevalent business practice of merely identifying and removing fraudulent reviews, as their study demonstrates that such an approach can result in persistent, long-term negative effects. This highlights the need for a robust statistical approach to protect against adversarial manipulation in matrix completion for discrete rating-scale data, which RDMC achieves by incorporating a robust loss function.
@@
Third, much of the literature does not investigate the effect of missing data mechanisms other than missing completely at random (MCAR) on matrix completion procedures, with some notable exceptions being Mao et al., (2019) for missing at random (MAR) as well as Choi & Yuan, (2024) and Xu et al., (2025) for missing not at random (MNAR). Yet missing values in recommender systems applications are not MCAR in practice (LeBlanc et al.,, 2024): users predominantly rate items that they have consumed or purchased and thus expected to like, intrinsically linking the probability of missingness to their preferences. Hence, we study the performance of the proposed method under realistic MNAR settings.
For transparency and reproducibility—key concerns in the current scientific landscape as discussed in LeBlanc et al., (2024)—we provide implementations of the methodology in package RMCLab (Alfons & Archimbaud,, 2025) for the statistical computing environment R (R Core Team,, 2025), which is available from https://CRAN.R-project.org/package=RMCLab. The main part of the code is thereby written in C++ to improve computational efficiency. Furthermore, replication files of all analyses will be made publicly available upon acceptance of this manuscript.
@@
The remainder of the paper is structured as follows. Section 2 introduces the proposed procedure RDMC for robust matrix completion with discrete rating-scale data. In Section 3, we introduce a simulation design that mimics recommender systems applications, and we investigate the performance of RDMC. Section 4 then presents two empirical case studies. The final Section 5 provides a concluding discussion.
@@ 2 A robust procedure for discrete matrix completion
We start by formulating the regularized optimization problem for matrix completion in Section 2.1 and presenting the corresponding algorithm in Section 2.2. Section 2.3 provides further algorithmic details, while Section 2.4 introduces different choices of robust loss functions. Finally, the selection of the regularization parameter is discussed in Section 2.5.
@@ 2.1 Problem formulation
Suppose that we observe an incomplete rating matrix $\bm{R}$ of $n$ rows (representing individuals providing the ratings) and $p$ columns (representing the rated items) with elements $R_{ij}\in\{1,2,\dots,K\}$. For practical reasons (see Section 2.3 for further discussion), we consider a column-centered matrix $\bm{X}$ with elements $X_{ij}\in\mathcal{C}_{j}=\{c_{1}^{j},\dots,c_{K}^{j}\}\subset\mathbb{R}$, where $c_{1}^{j}<\dots<c_{K}^{j}$ for $j=1,\dots,p$. Note that the values of the rating categories may vary between columns of $\bm{X}$. Hence, the assumption that the original rating matrix $\bm{R}$ takes values in $\{1,2,\dots,K\}$ is purely for simplicity and without loss of generality, and the proposed procedure extends in a straightforward manner to settings where even the number of rating categories may differ per column.
Let $\Omega$ denote the index set of observed entries in $\bm{X}$ and define the projection $P_{\Omega}$ to be the $(n\times p)$-dimensional matrix whose elements are given by
$\left(P_{\Omega}(\bm{X})\right)_{ij}=\begin{cases}X_{ij}&\text{ if }(i,j)\in\Omega,\\ 0&\text{ otherwise.}\end{cases}$
@@
In order to preserve the discrete nature of the ratings, one could add a discreteness constraint to (1) (cf. Huang et al.,, 2013). However, requiring the elements of $\bm{L}$ to be discrete may not yield a solution, since such a discreteness constraint and a rank constraint are unlikely to be fulfilled simultaneously. To resolve this issue, an ancillary continuous matrix $\bm{Z}$ can be introduced into problem (1) such that the discreteness constraint operates on $\bm{L}$ and the nuclear norm constraint on $\bm{Z}$, while ensuring that $\bm{L}$ and $\bm{Z}$ remain as close as possible. In addition, the squared Frobenius norm in (1) does not protect against corrupted observations (such as fake profiles in recommender systems). Indeed, this norm penalizes large errors quadratically, implying that corrupted observations with large errors may dominate the loss function and significantly influence the recommender system. For increased protection, we replace it with a robust loss function that puts less emphasis on large errors, thereby preventing them from disproportionately affecting the recommender system. More specifically, we replace it with a (pseudo-)norm $\|\bm{Y}\|_{\rho}=\sum_{i,j}\rho(Y_{ij})$ based on a robust loss function $\rho$ (cf. Tang & Guan,, 2020). Putting all of this together, we obtain the following optimization problem in augmented Lagrangian form:
@@
where $\langle\cdot,\cdot\rangle_{F}$ denotes the Frobenius inner product, $\bm{\Theta}$ is a multiplier adjusting for the discrepancy between $\bm{L}$ and $\bm{Z}$, and $\mu$ is an additional regularization parameter. It should be noted that (2) generalizes the optimization problem formulated in Huang et al., (2013) to a wider class of robust loss functions and by allowing different values of the rating categories between columns of the (column-centered) rating matrix.
@@ 2.2 Algorithm
$L_{ij}=\begin{cases}{\displaystyle\operatorname*{arg\,min}_{c_{k}\in\mathcal{C}_{j}}}\ \rho(c_{k}-X_{ij})+\frac{\mu}{2}(c_{k}-Z_{ij}+\frac{1}{\mu}\Theta_{ij})^{2}&\text{ for }(i,j)\in\Omega,\\ {\displaystyle\operatorname*{arg\,min}_{c_{k}\in\mathcal{C}_{j}}}\ (c_{k}-Z_{ij}+\frac{1}{\mu}\Theta_{ij})^{2}&\text{ for }(i,j)\notin\Omega.\end{cases}$ (4)
Third, update the discrepancy parameter $\bm{\Theta}\leftarrow\bm{\Theta}+\mu(\bm{L}-\bm{Z})$ and the regularization parameter $\mu\leftarrow\delta\mu$ with $\delta>1$. Hence, $\mu$ grows exponentially in the number of iterations to speed up convergence.
Pseudo-code for the algorithm is presented in Algorithm 1. Following Huang et al., (2013) and Tang & Guan, (2020), we initialize $\mu=0.1$ and set $\delta=1.05$. As convergence criterion, we take the relative change in the objective function from (2) falling below a given threshold, which we set to $\varepsilon_{\text{tol}}=10^{-4}$. Furthermore, we set the maximum number of iterations to $t_{\text{max}}=100$, as this sufficed for convergence in all our numerical experiments.
Algorithm 1 Robust matrix completion for discrete rating-scale data
Input Incomplete rating matrix $\bf R$, loss function $\rho(\cdot)$, regularization parameter $\lambda$, regularization parameter $\mu$, update factor $\delta$, convergence threshold $\varepsilon_{\text{tol}}$, maximum number of iterations $t_{\text{max}}$
   Output Complete rating matrix $\widehat{\bf R}$
[…]
9:  ${{\bf Z}^{(t)}\leftarrow\displaystyle\operatorname*{arg\,min}_{\bm{Z}}}{\textstyle\ \frac{1}{2}\|(\bm{L}^{(t-1)}+\frac{1}{\mu}\bm{\Theta}^{(t-1)})-\bm{Z}\|_{F}^{2}+\frac{\lambda}{\mu}\|\bm{Z}\|_{*}}$triangleright using Equation (3)
10:  ${\bf L}^{(t)}\leftarrow\displaystyle\operatorname*{arg\,min}_{{\bf L}}\|P_{\Omega}(\bm{X})-P_{\Omega}(\bm{L})\|_{\rho}+\frac{\mu}{2}\|\bm{L}-\bm{Z}^{(t)}+\frac{1}{\mu}\bm{\Theta}^{(t-1)}\|_{F}^{2}$      subject to $L_{ij}\in\mathcal{C}_{j}$triangleright using Equation (4)
11:  ${\bm{\Theta}}^{(t)}\leftarrow{\bm{\Theta}}^{(t-1)}+\mu({\bf L}^{(t)}-{\bf Z}^{(t)})$
12:  $\mu\leftarrow\delta\mu$
13:  $\text{Loss}^{(t)}\leftarrow\|P_{\Omega}({\bf X})-P_{\Omega}({\bf L}^{(t)})\|_{\rho}+\lambda\|{\bf Z}^{(t)}\|_{*}+\langle\bm{\Theta}^{(t)},{\bf L}^{(t)}-{\bf Z}^{(t)}\rangle_{F}+\frac{\mu}{2}\|{\bf L}^{(t)}-{\bf Z}^{(t)}\|_{F}^{2}$
14:  if $t>1$ then
15:   $\text{converged}\leftarrow|(\text{Loss}^{(t)}-\text{Loss}^{(t-1)})/\text{Loss}^{(t-1)}|\leq\varepsilon_{\text{tol}}$
16:  $t\leftarrow t+1$
[…]